Given a short run cost function as T C =1 /Q 3 (3 is under subscript ) - 2Q 2(2 is under subscript ) +60Q+100 , find the minimum value of AVC and MC



1
Expert's answer
2021-08-26T11:45:10-0400

"TC=\frac{1}{3}Q^3-2Q^2+60Q+100"

"TVC=\frac{1}{3}Q^3-2Q^2+60Q"


To find the minimum value of MC,

"MC=\frac{dC}{dQ}\:"

"MC=Q^2-4Q+60"

"\frac{dMC}{dQ}=2Q-4=0"

"Q=\frac{4}{2}"

"Q=2"


To find the minimum value of AVC,

"AVC=\frac{TVC}{Q}=\frac{\frac{1}{3}Q^3-2Q^2+60Q}{Q}"

"AVC=\frac{1}{3}\:Q^2-2Q+60"

"\frac{dAVC\:}{dQ}=\frac{2}{3}Q-2=0"

"\frac{2}{3}Q=2"

"Q=3"